# Dual thermal pseudo-critical features in a spin-1/2 Ising chain with twin-diamond geometry

### v2

@@ Abstract
We study the coupled twin-diamond chain, a decorated one-dimensional Ising model motivated by the magnetic structure of $\mathrm{Cu}_{2}(\mathrm{TeO}_{3})_{2}\mathrm{Br}_{2}$. By applying an exact mapping to an effective Ising chain, we obtain the full thermodynamic description of the system through a compact transfer-matrix formulation. The ground-state analysis reveals five distinct phases, including two frustrated sectors with extensive degeneracy. These frustrated regions give rise to characteristic entropy plateaus and separate the ordered phases in the zero-temperature diagram. At low temperatures the model exhibits peculiar sharp yet continuous variations of entropy, magnetization, and response functions, reflecting clear signatures of pseudo-transition behavior. The coupled twin-diamond chain thus provides an exactly solvable setting in which competing local configurations and internal frustration lead to pronounced dual pseudo-critical features in one dimension.
Pseudo-transition; Spin chain models; decorated Ising model.
@@
One-dimensional lattice models have recently attracted renewed attention because certain systems display thermodynamic anomalies that closely resemble finite-temperature phase transitions (galisova, ; strecka16, ; sm-souza16, ). These anomalies, known as pseudo-transitions, differ from genuine criticality even though they reproduce several of its characteristic signatures. First derivatives of the free energy, such as entropy and magnetization, exhibit steep but continuous variations near a pseudo-critical temperature, while second derivatives, including the specific heat and magnetic susceptibility, develop sharp yet finite peaks (sergio18, ). Numerical and analytical studies of decorated Ising and Ising-Heisenberg chains have further shown that these pseudo-transitions can mimic essential features of critical phenomena: the correlation length increases rapidly near the pseudo-critical temperature (Icarvalho, ), and several response functions follow power-law forms governed by a universal set of pseudo-critical exponents identified in many one-dimensional decorated models (Rojas2019, ).
@@
Peculiar decorated one-dimensional spin models provide a natural framework for these anomalies because their internal degrees of freedom generate competing low-energy sectors whose exchange of dominance at a characteristic temperature produces fully analytic yet remarkably sharp thermodynamic crossovers. This mechanism has been identified in several exactly solvable systems, including the Ising-XYZ diamond chain, where quasi-phases and pseudo-transitions were first clearly characterized (sergio18, ; Icarvalho, ), the coupled spin-electron double-tetrahedral chain, which exhibits frustration-enhanced entropy plateaus (galisova, ), and various Ising-Heisenberg ladders and tubes that display pseudo-critical behavior governed by universal exponents (strecka16, ; Rojas2019, ; Derzhko, ; dmd-potts, ; Krokhmalskii, ; Chapman, ). A unified interpretation based on the decoration-iteration mapping clarified that the anomalous crossover originates from temperature-dependent effective couplings and effective fields generated after integrating out the decorated sublattices (Rojas2019, ). Furthermore, subsequent analyses established that the residual entropy at zero-temperature phase boundaries serves as a reliable predictor for the appearance of pseudo-transitions in decorated chains (bjp20, ).
@@
Pseudo-transition behavior is not restricted to decorated chains. Several strictly one-dimensional models without auxiliary spins or internal bond structures also display sharp thermal crossovers when their parameters are tuned to generate competing low-energy configurations. A notable example is the extended Hubbard model at half filling in the atomic limit, where pronounced thermodynamic anomalies mimic those associated with quasi-phases (verissimo, ). A similar crossover between antiferromagnetic and charge-ordered configurations was reported in a minimal spin-pseudospin description of cuprate chains, accompanied by a strong but finite specific heat peak (katarina, ). Related effects occur in frustrated Potts-type models such as the q-state Potts-Zimm-Bragg formulation (Panov2021, ), in dilute Ising chains with impurity induced frustration (Yasinskaya, ), and in water-like lattice models incorporating van der Waals interactions and directional bonding (Braz25, ). These examples demonstrate that the essential ingredients behind pseudo-transitions, namely competing sectors with small energy gaps and enhanced degeneracies, can arise naturally even in purely one-dimensional systems without explicit decoration.
@@
Diamond chain geometries play a central role in several magnetic compounds, which explains why diamond chain models have been extensively studied in the theoretical literature (sm-souza16, ; strecka-dmd, 2020, 2004, ). A prominent example is the natural mineral azurite $\mathrm{Cu_{3}(CO_{3})_{2}(OH)_{2}}$(kikuchi03, ; kikuchi04, ), whose magnetic properties motivated detailed analyses of spin 1/2 Ising and Ising-Heisenberg diamond chains. Other compounds, such as the mixed valent iron diamond chain with single ion anisotropy (Sorolla, ), further support the idea that the diamond motif is an efficient source of frustration and anomalous thermodynamic behavior in one dimensional systems, among other similar structures(Derzhko, ; zad, ; Chakhmakhchyan, ). For these reasons, diamond chains remain key reference models for exploring the origin of entropy plateaus, degeneracy driven anomalies, and pseudo-critical behavior in one dimension.
@@
Beyond simple diamond chains, an increasing number of materials display coupled diamond like building blocks, often arranged in layers or extended structural networks. A representative case is the oxohalide $\mathrm{Cu}_{2}(\mathrm{TeO}_{3})_{2}\mathrm{Br}_{2}$, where $\mathrm{Cu}^{2+}$ ions form diamond shaped units that are connected along a crystallographic direction (Uematsu, 2007). Similar motifs appear in more complex families such as the $\mathrm{CuO\text{\textendash}CuCl_{2}\text{\textendash}SeO_{2}}$ series (Kakarla, ), as well as in the oxohalides $\mathrm{Cu}_{3}(\mathrm{SeO}_{3})_{2}\mathrm{Cl}_{2}$ and $\mathrm{Cu}_{3}(\mathrm{TeO}_{3})_{2}\mathrm{Br}_{2}$ (Rocquefelte, ), and the insulating spin-gap compound $\mathrm{CdCu_{2}(SeO_{3})_{2}Cl_{2}}$ (Murtazoev, ). Such structural features strongly motivate the study of theoretical models that incorporate more than one diamond unit per cell. The coupled twin diamond chain examined in the present work represents the minimal extension capable of capturing this additional connectivity and allows us to investigate how the interaction between two diamond units within each cell can produce multiple frustrated sectors and distinct pseudo-critical features in a single exactly solvable framework.
@@
In the present work, the coupled twin-diamond chain is therefore introduced as a minimal theoretical model inspired by such diamond-based magnetic structures(Uematsu, 2007), rather than as a quantitative description of any specific compound. The central novelty of the coupled twin-diamond chain lies in the fact that its minimal unit cell supports more than one frustrated low-energy manifold with different extensive degeneracies. As a result, the system exhibits two well-separated pseudo-critical temperature scales associated with successive entropy-driven crossovers between competing configurations. This provides a clean exactly solvable example in which dual pseudo-transition features arise within a single one-dimensional decorated Ising geometry, rather than from tuning parameters between distinct models. We further show that these features remain robust even when the two spin sublattices respond differently to an external magnetic field.
@@
This work is organized as follows. Sect. II introduces the Hamiltonian, outlines the transfer-matrix solution, and presents the zero-temperature phase diagram together with the corresponding pseudo-transition features. Sect. III examines the thermodynamic behavior in the anomalous region where these pseudo-transitions emerge. Finally, Sec. IV summarizes the main conclusions.
@@ II.1 Hamiltonian and transfer matrix
$y=$ $2\sqrt{Q}\cos\left(\tfrac{R}{3\sqrt{Q^{3}}}\right)+\frac{a_{2}}{6},$ (19)
@@ III.2 Phase boundary
and these do not exhibit any enhancement of the boundary residual entropy. It is worth noting that both of these boundaries produce clear pseudo-transition signatures at low temperature. In Fig. 2 they manifest as sharp entropic ridges that separate the ordered and frustrated regions in the $\mathrm{FI}-\mathrm{FR}_{1}$ case, and similarly as a distinct ridge between the two frustrated states in the $\mathrm{FR}_{1}-\mathrm{FR}_{2}$ case.
@@ III.3 Nearly block-diagonal form of the transfer matrix
Since the transfer matrix $\mathbf{V}$ given in Eq. (2) is irreducible, the Non-existence of phase transition theorem(cuesta, ) guarantees the absence of true phase transitions. Nevertheless, $\mathbf{V}$ may become nearly block diagonal in certain parameter regimes. In this case, the approach proposed in Ref. (gen-conds, ) provides a useful framework to analyze pseudo-critical behavior from a general perspective, without this tool it would become a cumbersome task find the pseudo-critical temperature. Following this approach, we reorganize the transfer matrix into a nearly block-diagonal form in order to identify the competing thermodynamic sectors responsible for the observed pseudo-critical temperatures.
@@ III.3.1 Pair-sector decomposition
In terms of the ground-state patterns discussed above, the $\mathrm{FR}_{2}$ and FI manifolds are predominantly supported on parallel sector, whereas the $\mathrm{FR}_{1}$ manifold is supported on antiparallel sector. Consequently, the $\mathrm{FI}-\mathrm{FR}_{1}$ and $\mathrm{FR}_{1}-\mathrm{FR}_{2}$ pseudo-transitions can be viewed as successive competitions between these two pair sectors.
@@ III.3.2 Nearly block-diagonal structure of the transfer matrix
Although $\mathbf{V}$ is irreducible, the inter-sector couplings $\mathbf{C}$ and $\mathbf{D}$ can become parametrically small in the low-temperature regime. In this sense, the transfer matrix is nearly block diagonal: the dominant statistical weight is carried by one sector at a time, and pseudo-critical behavior emerges when the leading eigenvalues associated with different sectors become nearly degenerate. This nearly reducible structure provides the spectral mechanism underlying the sharp yet analytic crossovers discussed in the main text.
@@
III.3.3 Pseudo-critical temperature
According to the criterion discussed in Ref. (gen-conds, ), pseudo-critical behavior arises when the transfer matrix acquires a nearly block-diagonal structure, such that the leading eigenvalues associated with competing blocks become comparable. In the present case, this leads to two possible pseudo-critical conditions, depending on which eigenvalue of the block $\mathbf{A}$ is dominant:
@@ III.3.3 Pseudo-critical temperature
These relations provide a transparent interpretation of the two pseudo-critical temperature scales observed in the thermodynamic analysis. It is worth emphasizing that they follow directly from the criterion of Ref. (gen-conds, ). Although simple in form, the resulting conditions involve transcendental equations.
Alternatively, the pseudo-critical temperature $T_{p}$ can also be determined along any fixed parameter cut in the low-temperature regime. For instance at fixed $J_{0}$, $J_{1}$, $J_{2}$, while varying $B$, follows from the crossing of the low temperature free energy branches of the two competing phases. Approximating each free energy by $f_{A}(T)\simeq\varepsilon_{A}-Ts_{A}$, the condition $f_{A}(T_{\mathrm{p}})=f_{B}(T_{\mathrm{p}})$ yields the standard estimate of pseudo-critical temperature
$k_{\mathrm{B}}T_{\mathrm{p}}\approx\frac{E_{B}-E_{A}}{\mathcal{S}_{B}-\mathcal{S}_{A}}.$ (35)
@@
which accurately reproduces the low-temperature behavior of $T_{p_{1}}$ obtained in Eq. (33). In a similarly manner, the pseudo-critical temperature associated with the $\mathrm{FR}_{1}$ and $\mathrm{FR}_{2}$ competition is estimated as
@@
In Fig. 3 is shown, how the low-temperature entropy $S(T)$ landscape changes in the $(J_{2}/J_{1},\,\mu_{{\rm B}}g_{0}B/J_{1})$ plane, when the two sublattices have $g_{0}\neq g_{1}$. In panel (a), where $g_{1}/g_{0}>1$, the frustrated boundaries that appear as vertical lines in the symmetric case tilt into straight lines with a positive slope. In panel (b), where $g_{1}/g_{0}<1$, the same boundaries acquire a negative slope, producing a opposite deformation of the diagram.
@@
Importantly, the deformation of these boundaries does not suppress the thermal anomalies associated with the frustrated manifold. The tilted ridges that separate FI from $\mathrm{FR}_{1}$, and $\mathrm{FR}_{1}$ from $\mathrm{FR}_{2}$ continue to generate clear pseudo-transition signatures at low temperature, reflected in the sharp entropic gradients that survive in both panels.
Figure 3: Low-temperature phase diagram of the CTDC for $g_{0}\neq g_{1}$ under a uniform external field $B$, assuming fixe $J_{0}/J_{1}=-0.1$. (a) Dimer spins have weaker Zeeman response, $g_{1}/g_{0}=0.7$. (b) Dimer spins have stronger Zeeman response, $g_{1}/g_{0}=1.3$. The entropy plot at $k_{{\rm B}}T/J_{1}=0.05$ displays the corresponding finite-temperature structure.
IV Thermodynamics and pseudo-transition

### v3

# Dual thermal pseudocritical features in a spin-1/2 Ising chain with twin-diamond geometry

@@ Abstract
We study the coupled twin-diamond chain, a decorated one-dimensional Ising model motivated by the magnetic structure of $\mathrm{Cu}_{2}(\mathrm{TeO}_{3})_{2}\mathrm{Br}_{2}$. By applying an exact mapping to an effective Ising chain, we obtain the full thermodynamic description of the system through a compact transfer-matrix formulation. The ground-state analysis reveals five distinct phases, including two frustrated sectors with extensive degeneracy. These frustrated regions give rise to characteristic entropy plateaus and separate the ordered phases in the zero-temperature diagram. At low temperatures the model exhibits peculiar sharp yet continuous variations of entropy, magnetization, and response functions, reflecting clear signatures of pseudotransition behavior. The coupled twin-diamond chain thus provides an exactly solvable setting in which competing local configurations and internal frustration lead to pronounced dual pseudocritical features in one dimension.
Pseudotransition; Spin chain models; decorated Ising model.
@@
One-dimensional lattice models have recently attracted renewed attention because certain systems display thermodynamic anomalies that closely resemble finite-temperature phase transitions (galisova, ; strecka16, ; sm-souza16, ). These anomalies, known as pseudotransitions, differ from genuine criticality even though they reproduce several of its characteristic signatures. First derivatives of the free energy, such as entropy and magnetization, exhibit steep but continuous variations near a pseudocritical temperature, while second derivatives, including the specific heat and magnetic susceptibility, develop sharp yet finite peaks (sergio18, ). Numerical and analytical studies of decorated Ising and Ising-Heisenberg chains have further shown that these pseudotransitions can mimic essential features of critical phenomena: the correlation length increases rapidly near the pseudocritical temperature (Icarvalho, ), and several response functions follow power-law forms governed by a universal set of pseudocritical exponents identified in many one-dimensional decorated models (Rojas2019, ).
@@
Peculiar decorated one-dimensional spin models provide a natural framework for these anomalies because their internal degrees of freedom generate competing low-energy sectors whose exchange of dominance at a characteristic temperature produces fully analytic yet remarkably sharp thermodynamic crossovers. This mechanism has been identified in several exactly solvable systems, including the Ising-XYZ diamond chain, where quasi-phases and pseudotransitions were first clearly characterized (sergio18, ; Icarvalho, ), the coupled spin-electron double-tetrahedral chain, which exhibits frustration-enhanced entropy plateaus (galisova, ), and various Ising-Heisenberg ladders and tubes that display pseudocritical behavior governed by universal exponents (strecka16, ; Rojas2019, ; Derzhko, ; dmd-potts, ; Krokhmalskii, ; Chapman, ). A unified interpretation based on the decoration-iteration mapping clarified that the anomalous crossover originates from temperature-dependent effective couplings and effective fields generated after integrating out the decorated sublattices (Rojas2019, ). Furthermore, subsequent analyses established that the residual entropy at zero-temperature phase boundaries serves as a reliable predictor for the appearance of pseudotransitions in decorated chains (bjp20, ).
@@
Pseudotransition behavior is not restricted to decorated chains. Several strictly one-dimensional models without auxiliary spins or internal bond structures also display sharp thermal crossovers when their parameters are tuned to generate competing low-energy configurations. A notable example is the extended Hubbard model at half filling in the atomic limit, where pronounced thermodynamic anomalies mimic those associated with quasi-phases (verissimo, ). A similar crossover between antiferromagnetic and charge-ordered configurations was reported in a minimal spin-pseudospin description of cuprate chains, accompanied by a strong but finite specific heat peak (katarina, ). Related effects occur in frustrated Potts-type models such as the q-state Potts-Zimm-Bragg formulation (Panov2021, ), in dilute Ising chains with impurity induced frustration (Yasinskaya, ), and in water-like lattice models incorporating van der Waals interactions and directional bonding (Braz25, ). These examples demonstrate that the essential ingredients behind pseudotransitions, namely competing sectors with small energy gaps and enhanced degeneracies, can arise naturally even in purely one-dimensional systems without explicit decoration.
@@
Diamond chain geometries play a central role in several magnetic compounds, which explains why diamond chain models have been extensively studied in the theoretical literature (sm-souza16, ; strecka-dmd, 2020, 2004, ). A prominent example is the natural mineral azurite $\mathrm{Cu_{3}(CO_{3})_{2}(OH)_{2}}$(kikuchi03, ; kikuchi04, ), whose magnetic properties motivated detailed analyses of spin 1/2 Ising and Ising-Heisenberg diamond chains. Other compounds, such as the mixed valent iron diamond chain with single ion anisotropy (Sorolla, ), further support the idea that the diamond motif is an efficient source of frustration and anomalous thermodynamic behavior in one dimensional systems, among other similar structures(Derzhko, ; zad, ; Chakhmakhchyan, ). For these reasons, diamond chains remain key reference models for exploring the origin of entropy plateaus, degeneracy driven anomalies, and pseudocritical behavior in one dimension.
@@
Beyond simple diamond chains, an increasing number of materials display coupled diamond like building blocks, often arranged in layers or extended structural networks. A representative case is the oxohalide $\mathrm{Cu}_{2}(\mathrm{TeO}_{3})_{2}\mathrm{Br}_{2}$, where $\mathrm{Cu}^{2+}$ ions form diamond shaped units that are connected along a crystallographic direction (Uematsu, 2007). Similar motifs appear in more complex families such as the $\mathrm{CuO\text{\textendash}CuCl_{2}\text{\textendash}SeO_{2}}$ series (Kakarla, ), as well as in the oxohalides $\mathrm{Cu}_{3}(\mathrm{SeO}_{3})_{2}\mathrm{Cl}_{2}$ and $\mathrm{Cu}_{3}(\mathrm{TeO}_{3})_{2}\mathrm{Br}_{2}$ (Rocquefelte, ), and the insulating spin-gap compound $\mathrm{CdCu_{2}(SeO_{3})_{2}Cl_{2}}$ (Murtazoev, ). Such structural features strongly motivate the study of theoretical models that incorporate more than one diamond unit per cell. The coupled twin diamond chain examined in the present work represents the minimal extension capable of capturing this additional connectivity and allows us to investigate how the interaction between two diamond units within each cell can produce multiple frustrated sectors and distinct pseudocritical features in a single exactly solvable framework.
@@
In the present work, the coupled twin-diamond chain is therefore introduced as a minimal theoretical model inspired by such diamond-based magnetic structures(Uematsu, 2007), rather than as a quantitative description of any specific compound. The central novelty of the coupled twin-diamond chain lies in the fact that its minimal unit cell supports more than one frustrated low-energy manifold with different extensive degeneracies. As a result, the system exhibits two well-separated pseudocritical temperature scales associated with successive entropy-driven crossovers between competing configurations. This provides a clean exactly solvable example in which dual pseudotransition features arise within a single one-dimensional decorated Ising geometry, rather than from tuning parameters between distinct models. We further show that these features remain robust even when the two spin sublattices respond differently to an external magnetic field.
@@
This work is organized as follows. Sect. II introduces the Hamiltonian, outlines the transfer-matrix solution, and presents the zero-temperature phase diagram together with the corresponding pseudotransition features. Sect. III examines the thermodynamic behavior in the anomalous region where these pseudotransitions emerge. Finally, Sec. IV summarizes the main conclusions.
@@ II.1 Hamiltonian and transfer matrix
$y=$ $2\sqrt{Q}\cos\left(\tfrac{1}{3}\arccos(\tfrac{R}{\sqrt{Q^{3}}})\right)+\frac{a_{2}}{6},$ (19)
@@ III.2 Phase boundary
and these do not exhibit any enhancement of the boundary residual entropy. It is worth noting that both of these boundaries produce clear pseudotransition signatures at low temperature. In Fig. 2 they manifest as sharp entropic ridges that separate the ordered and frustrated regions in the $\mathrm{FI}-\mathrm{FR}_{1}$ case, and similarly as a distinct ridge between the two frustrated states in the $\mathrm{FR}_{1}-\mathrm{FR}_{2}$ case.
@@ III.3 Nearly block-diagonal form of the transfer matrix
Since the transfer matrix $\mathbf{V}$ given in Eq. (2) is irreducible, the Non-existence of phase transition theorem(cuesta, ) guarantees the absence of true phase transitions. Nevertheless, $\mathbf{V}$ may become nearly block diagonal in certain parameter regimes. In this case, the approach proposed in Ref. (gen-conds, ) provides a useful framework to analyze pseudocritical behavior from a general perspective, without this tool it would become a cumbersome task find the pseudocritical temperature. Following this approach, we reorganize the transfer matrix into a nearly block-diagonal form in order to identify the competing thermodynamic sectors responsible for the observed pseudocritical temperatures.
@@ III.3.1 Pair-sector decomposition
In terms of the ground-state patterns discussed above, the $\mathrm{FR}_{2}$ and FI manifolds are predominantly supported on parallel sector, whereas the $\mathrm{FR}_{1}$ manifold is supported on antiparallel sector. Consequently, the $\mathrm{FI}-\mathrm{FR}_{1}$ and $\mathrm{FR}_{1}-\mathrm{FR}_{2}$ pseudotransitions can be viewed as successive competitions between these two pair sectors.
@@ III.3.2 Nearly block-diagonal structure of the transfer matrix
Although $\mathbf{V}$ is irreducible, the inter-sector couplings $\mathbf{C}$ and $\mathbf{D}$ can become parametrically small in the low-temperature regime. In this sense, the transfer matrix is nearly block diagonal: the dominant statistical weight is carried by one sector at a time, and pseudocritical behavior emerges when the leading eigenvalues associated with different sectors become nearly degenerate. This nearly reducible structure provides the spectral mechanism underlying the sharp yet analytic crossovers discussed in the main text.
@@
III.3.3 Pseudocritical temperature
According to the criterion discussed in Ref. (gen-conds, ), pseudocritical behavior arises when the transfer matrix acquires a nearly block-diagonal structure, such that the leading eigenvalues associated with competing blocks become comparable. In the present case, this leads to two possible pseudocritical conditions, depending on which eigenvalue of the block $\mathbf{A}$ is dominant:
@@ III.3.3 Pseudocritical temperature
These relations provide a transparent interpretation of the two pseudocritical temperature scales observed in the thermodynamic analysis. It is worth emphasizing that they follow directly from the criterion of Ref. (gen-conds, ). Although simple in form, the resulting conditions involve transcendental equations.
Alternatively, the pseudocritical temperature $T_{p}$ can also be determined along any fixed parameter cut in the low-temperature regime. For instance at fixed $J_{0}$, $J_{1}$, $J_{2}$, while varying $B$, follows from the crossing of the low temperature free energy branches of the two competing phases. Approximating each free energy by $f_{A}(T)\simeq\varepsilon_{A}-k_{\mathrm{B}}Ts_{A}$, the condition $f_{A}(T_{\mathrm{p}})=f_{B}(T_{\mathrm{p}})$ yields the standard estimate of pseudocritical temperature
$k_{\mathrm{B}}T_{\mathrm{p}}\approx\frac{\varepsilon_{B}-\varepsilon_{A}}{\mathcal{S}_{B}-\mathcal{S}_{A}}.$ (35)
@@
which accurately reproduces the low-temperature behavior of $T_{p_{1}}$ obtained in Eq. (33). In a similarly manner, the pseudocritical temperature associated with the $\mathrm{FR}_{1}$ and $\mathrm{FR}_{2}$ competition is estimated as
@@
In Fig. 3 is shown, how the low-temperature entropy $S(T)$ landscape changes in the $(J_{2}/J_{1},\,\mu_{{\rm B}}g_{0}B/J_{1})$ plane, when the two sublattices have $g_{0}\neq g_{1}$. In panel (left), where $g_{1}/g_{0}>1$, the frustrated boundaries that appear as vertical lines in the symmetric case tilt into straight lines with a positive slope. In panel (right), where $g_{1}/g_{0}<1$, the same boundaries acquire a negative slope, producing a opposite deformation of the diagram.
@@
Importantly, the deformation of these boundaries does not suppress the thermal anomalies associated with the frustrated manifold. The tilted ridges that separate FI from $\mathrm{FR}_{1}$, and $\mathrm{FR}_{1}$ from $\mathrm{FR}_{2}$ continue to generate clear pseudotransition signatures at low temperature, reflected in the sharp entropic gradients that survive in both panels.
Figure 3: Low-temperature phase diagram of the CTDC for $g_{0}\neq g_{1}$ under a uniform external field $B$, assuming fixe $J_{0}/J_{1}=-0.1$. (left) Dimer spins have weaker Zeeman response, $g_{1}/g_{0}=0.7$. (right) Dimer spins have stronger Zeeman response, $g_{1}/g_{0}=1.3$. The entropy plot at $k_{{\rm B}}T/J_{1}=0.05$ displays the corresponding finite-temperature structure.
IV Thermodynamics and pseudotransition